# Optimal low-rank approximations of posteriors for linear Gaussian inverse problems on Hilbert spaces

### v1

@@ B.1 Proofs of Section 3
We now prove the stated properties of the eigenvalues and eigenvectors. Recall that by (2), $H\in\mathcal{B}_{00,n}(\mathcal{H})$ is self-adjoint and non-negative. Hence $\mathcal{C}_{\textup{pr}}^{1/2}H\mathcal{C}_{\textup{pr}}^{1/2}=(\mathcal{C}_{\textup{pr}}^{1/2}H^{1/2})(\mathcal{C}_{\textup{pr}}^{1/2}H^{1/2})^{*}$ is also self-adjoint and non-negative, which implies that $(\tfrac{-\lambda_{i}}{1+\lambda_{i}})_{i\in\mathbb{N}}\subset\ell^{2}((-1,0])$, and thus that $(\lambda_{i})_{i\in\mathbb{N}}\in\ell^{2}((-1,0])$. Since $\mathcal{C}_{\textup{pr}}$ is injective on $\mathcal{H}$, it follows that $\operatorname{rank}\left({\mathcal{C}_{\textup{pr}}^{1/2}H\mathcal{C}_{\textup{pr}}^{1/2}}\right)=\operatorname{rank}\left({H}\right)$, whence $(\tfrac{-\lambda_{i}}{1+\lambda_{i}})_{i\in\mathbb{N}}$ contains exactly $\operatorname{rank}\left({H}\right)\leq n$ many nonzero entries. It follows directly from (10a), (10b) and the fact that $\lambda_{i}\not=0$ for $i\leq\operatorname{rank}\left({H}\right)$, that $w_{i}\in\operatorname{ran}{\mathcal{C}_{\textup{pr}}^{1/2}H\mathcal{C}_{\textup{pr}}^{1/2}}=\operatorname{ran}{\mathcal{C}_{\textup{pr}}^{1/2}}\subset\operatorname{ran}{\mathcal{C}_{\textup{pr}}^{1/2}}$ and $v_{i}\in\operatorname{ran}{\mathcal{C}_{\textup{pos}}^{1/2}H\mathcal{C}_{\textup{pos}}^{1/2}}\subset\operatorname{ran}{\mathcal{C}_{\textup{pos}}^{1/2}}=\operatorname{ran}{\mathcal{C}_{\textup{pr}}^{1/2}}$ for $i\leq\operatorname{rank}\left({H}\right)$. By Lemma A.2, we can extend $(w_{i})_{i=1}^{\operatorname{rank}\left({H}\right)}$ to an ONB $(w_{i}^{\prime})_{i}$ of $\mathcal{H}$ with $(w_{i}^{\prime})_{i}\subset\operatorname{ran}{\mathcal{C}_{\textup{pr}}^{1/2}}$ and $w_{i}^{\prime}=w_{i}$ for $i\leq\operatorname{rank}\left({H}\right)$. We now replace $w_{i}$ by $w_{i}^{\prime}$ and $v_{i}$ by $\mathcal{C}_{\textup{pos}}^{-1/2}\mathcal{C}_{\textup{pr}}^{1/2}w_{i}^{\prime}$ for $i>\operatorname{rank}\left({H}\right)$. The equations (10a) and (10b) and $v_{i}=\sqrt{1+\lambda_{i}}\mathcal{C}_{\textup{pos}}^{-1/2}\mathcal{C}_{\textup{pr}}^{1/2}w_{i}$ for all $i$ remain to hold, and we now have $w_{i},v_{i}\in\operatorname{ran}{\mathcal{C}_{\textup{pr}}^{1/2}}$ for all $i$.
By (i) of Lemma 3.3 and the fact that $(w_{i})_{i}$ lies in the Cameron–Martin space, it follows that
@@
Finally, we prove (11). Multiplying the equation (10c) from the left by $\mathcal{C}_{\textup{pos}}^{1/2}$ implies that $\mathcal{C}_{\textup{pos}}\mathcal{C}_{\textup{pr}}^{-1/2}w_{i}=(1+\lambda_{i})\mathcal{C}_{\textup{pr}}^{1/2}w_{i}=(1+\lambda_{i})\mathcal{C}_{\textup{pr}}(\mathcal{C}_{\textup{pr}}^{-1/2}w_{i})$. Taking the inner product of both sides of the last equation with $\mathcal{C}_{\textup{pr}}^{-1/2}w_{i}$, we obtain $\langle\mathcal{C}_{\textup{pos}}\mathcal{C}_{\textup{pr}}^{-1/2}w_{i},\mathcal{C}_{\textup{pr}}^{-1/2}w_{i}\rangle=(1+\lambda_{i})\langle\mathcal{C}_{\textup{pr}}\mathcal{C}_{\textup{pr}}^{-1/2}w_{i},\mathcal{C}_{\textup{pr}}^{-1/2}w_{i}\rangle$. By Lemma A.22, $\textup{Var}_{X\sim\mu_{\textup{pos}}}(\langle X,z\rangle)=\langle\mathcal{C}_{\textup{pos}}z,z\rangle$ and $\textup{Var}_{X\sim\mu_{\textup{pr}}}(\langle X,z\rangle)=\langle\mathcal{C}_{\textup{pr}}z,z\rangle$ for any $z\in\mathcal{H}$. Thus we obtain (11). ∎
@@
We use the expressions for the KL and Rényi divergence of [22, Theorem 14, Theorem 15]. While they are stated for infinite-dimensional Hilbert spaces only, it is noted in [23, after Theorem 3] that these expressions also hold for finite-dimensional Hilbert spaces. The statements in the theorem now follow immediately from the expressions in [22, Theorem 14, Theorem 15] after noting that, since $(\mathcal{C}_{1}^{-1/2}\mathcal{C}_{2}^{1/2})^{*}=\mathcal{C}_{2}^{1/2}\mathcal{C}_{1}^{-1/2}$ on $\operatorname{ran}{\mathcal{C}_{1}^{1/2}}$ by Lemma A.13(i), we have for $S\coloneqq-R(\mathcal{C}_{2}\|\mathcal{C}_{1})\in L_{2}(\mathcal{H})_{\mathbb{R}}$, where $R(\cdot\|\cdot)$ is defined in (9), that
@@ B.2 Proofs of Section 4
Let $(x_{i})_{i}\in\ell^{2}((-1,\infty))$. For $N$ large enough, its tail $(x_{i})_{i>N}$ lies in $(-\varepsilon,\varepsilon)$, so that the inequality above implies $\sum_{i>N}f(x_{i})\leq\tfrac{1}{2}M\sum_{i>N}x_{i}^{2}\leq\tfrac{1}{2}M\lVert x\rVert_{\ell^{2}}^{2}$. For $\mathcal{C}_{1}$, $\mathcal{C}_{2}\in\mathcal{E}$ we have $R(\mathcal{C}_{2}\|\mathcal{C}_{1})\in L_{2}(\mathcal{H})$ and its eigenvalue sequence is square summable. Hence $\mathcal{L}_{f}<\infty$ by the definition of $\mathcal{L}_{f}$ in (15a). This proves item (i). For item (ii), we note $f(\alpha(0))=f(0)=0$. Furthermore, we compute $\alpha^{\prime}(x)=-(1+x)^{-2}>0$ and, by the fact that $f\in\mathscr{F}$, $x(f\circ\alpha)^{\prime}(x)=\frac{1}{1+x}\frac{-x}{1+x}f^{\prime}(\frac{1}{1+x})>0$ for $x\not=0$. By the assumption on $f$, $\lim_{x\rightarrow\infty}f(\alpha(x))=\lim_{x\rightarrow-1}f(x)=\infty$. Finally, $\alpha$ is smooth, so $\alpha$ and $\alpha^{\prime}$ are Lipschitz at 0. Therefore, $f^{\prime}\circ\alpha$ is Lipschitz at 0 as the composition Lipschitz functions at 0 and $(f\circ\alpha)^{\prime}=(f^{\prime}\circ\alpha)\alpha^{\prime}$ is Lipschitz at 0 as the product of two Lipschitz functions at 0. ∎
@@
Hence $xf^{\prime}_{\textup{Ren},\rho}(x)>0$ for $x\not=0$. Furthermore, $f_{\textup{KL}}(0)=0=f_{\textup{Ren},\rho}(0)$ and $\lim_{x\rightarrow\infty}f_{\textup{KL}}(x)=\infty=\lim_{x\rightarrow\infty}f_{\textup{Ren},\rho}(x)$. So $f_{\textup{KL}},f_{\textup{Ren},\rho}\in\mathscr{F}$.
The first equations in items (i) and (ii) follows from Theorem 3.8. With $(\lambda_{i})_{i\in\mathbb{N}}$ the eigenvalues of $R(\mathcal{C}_{2}\|\mathcal{C}_{1})\in L_{2}(\mathcal{H})$, it holds by (13) that
@@
By the spectral mapping theorem, see e.g. [26, Theorem VII.1(e)] for a version that does not assume that $\mathcal{H}$ is defined over the complex field $\mathbb{C}$, the eigenvalues of $I+R(\mathcal{C}_{2}\|\mathcal{C}_{1})$ are $(1+\lambda_{i})_{i}$, and the eigenvalues of $A_{\rho}\coloneqq\bigl{(}I+R(\mathcal{C}_{2}\|\mathcal{C}_{1})\bigr{)}^{\rho-1}\bigl{(}\rho I+(1-\rho)(I+R(\mathcal{C}_{2}\|\mathcal{C}_{1}))\bigr{)}$ are $(\gamma_{i})_{i}$, with $\gamma_{i}\coloneqq(1+\lambda_{i})^{\rho-1}(\rho+(1-\rho)(1+\lambda_{i}))$, $i\in\mathbb{N}$. The eigenvalues of $A_{\rho}-I$ are then $(\gamma_{i}-1)_{i}$ and by (12),
$\det(A_{\rho})=\det(I+(A_{\rho}-I))=\prod_{i}(1+(\gamma_{i}-1))=\exp(\sum_{i}\log(\gamma_{i}))=\exp\left(2\rho(1-\rho)\sum_{i}f_{\textup{Ren},\rho}(\lambda_{i})\right).$
This shows item (ii) holds. Item (iii) now follows directly from items (i) and (ii) and Lemma 4.1(ii). ∎

### v3

@@ B.1 Proofs of Section 3
We now prove the stated properties of the eigenvalues and eigenvectors. Recall that by (2), $H\in\mathcal{B}_{00,n}(\mathcal{H})$ is self-adjoint and non-negative. Hence $\mathcal{C}_{\textup{pr}}^{1/2}H\mathcal{C}_{\textup{pr}}^{1/2}=(\mathcal{C}_{\textup{pr}}^{1/2}H^{1/2})(\mathcal{C}_{\textup{pr}}^{1/2}H^{1/2})^{*}$ is also self-adjoint and non-negative, which implies that $(\tfrac{-\lambda_{i}}{1+\lambda_{i}})_{i\in\mathbb{N}}\subset\ell^{2}((-1,0])$, and thus that $(\lambda_{i})_{i\in\mathbb{N}}\in\ell^{2}((-1,0])$. We thus may order $(\lambda_{i})_{i}$ in a nondecreasing manner. Since $\mathcal{C}_{\textup{pr}}$ is injective on $\mathcal{H}$, it follows by applying Lemma A.8 twice with $A\leftarrow\mathcal{C}_{\textup{pr}}^{1/2}H^{1/2}$ and $A\leftarrow H^{1/2}$ that
$\operatorname{rank}\left({\mathcal{C}_{\textup{pr}}^{1/2}H\mathcal{C}_{\textup{pr}}^{1/2}}\right)=\operatorname{rank}\left({\mathcal{C}_{\textup{pr}}^{1/2}H^{1/2}(\mathcal{C}_{\textup{pr}}^{1/2}H^{1/2})^{*}}\right)=\operatorname{rank}\left({\mathcal{C}_{\textup{pr}}^{1/2}H^{1/2}}\right)=\operatorname{rank}\left({H^{1/2}}\right)=\operatorname{rank}\left({H}\right).$
Therefore, $(\tfrac{-\lambda_{i}}{1+\lambda_{i}})_{i\in\mathbb{N}}$ contains exactly $\operatorname{rank}\left({H}\right)\leq n$ many nonzero entries. It follows directly from (10a), (10b) and the fact that $\lambda_{i}\not=0$ for $i\leq\operatorname{rank}\left({H}\right)$, that $w_{i}\in\operatorname{ran}{\mathcal{C}_{\textup{pr}}^{1/2}H\mathcal{C}_{\textup{pr}}^{1/2}}=\operatorname{ran}{\mathcal{C}_{\textup{pr}}^{1/2}}\subset\operatorname{ran}{\mathcal{C}_{\textup{pr}}^{1/2}}$ and $v_{i}\in\operatorname{ran}{\mathcal{C}_{\textup{pos}}^{1/2}H\mathcal{C}_{\textup{pos}}^{1/2}}\subset\operatorname{ran}{\mathcal{C}_{\textup{pos}}^{1/2}}=\operatorname{ran}{\mathcal{C}_{\textup{pr}}^{1/2}}$ for $i\leq\operatorname{rank}\left({H}\right)$. By Lemma A.1, we can extend $(w_{i})_{i=1}^{\operatorname{rank}\left({H}\right)}$ to an ONB $(w_{i}^{\prime})_{i}$ of $\mathcal{H}$ with $(w_{i}^{\prime})_{i}\subset\operatorname{ran}{\mathcal{C}_{\textup{pr}}^{1/2}}$ and $w_{i}^{\prime}=w_{i}$ for $i\leq\operatorname{rank}\left({H}\right)$. We now replace $w_{i}$ by $w_{i}^{\prime}$ and $v_{i}$ by $\mathcal{C}_{\textup{pos}}^{-1/2}\mathcal{C}_{\textup{pr}}^{1/2}w_{i}^{\prime}$ for $i>\operatorname{rank}\left({H}\right)$. The equations (10a) and (10b) and $v_{i}=\sqrt{1+\lambda_{i}}\mathcal{C}_{\textup{pos}}^{-1/2}\mathcal{C}_{\textup{pr}}^{1/2}w_{i}$ for all $i$ remain to hold, and we now have $w_{i},v_{i}\in\operatorname{ran}{\mathcal{C}_{\textup{pr}}^{1/2}}$ for all $i$.
By item (i) of Lemma 3.4 and the fact that $(w_{i})_{i}$ lies in the Cameron–Martin space, it follows that
@@
We use the expressions for the KL and Rényi divergence of [23, Theorem 14, Theorem 15]. While they are stated for infinite-dimensional Hilbert spaces only, it is noted in [24] that these expressions also hold for finite-dimensional Hilbert spaces; see the remarks after [24, Theorem 3]. By Lemma A.19(i), $(\mathcal{C}_{1}^{-1/2}\mathcal{C}_{2}^{1/2})^{*}=\mathcal{C}_{2}^{1/2}\mathcal{C}_{1}^{-1/2}$ on $\operatorname{ran}{\mathcal{C}_{1}^{1/2}}$. The statements in the theorem now follow immediately from the expressions in [23, Theorem 14, Theorem 15], because for $S\coloneqq-R(\mathcal{C}_{2}\|\mathcal{C}_{1})\in L_{2}(\mathcal{H})_{\mathbb{R}}$, where $R(\cdot\|\cdot)$ is defined in (9), we have
@@ B.2 Proofs of Section 4
Let $(x_{i})_{i}\in\ell^{2}((-1,\infty))$. For $N$ large enough, its tail $(x_{i})_{i>N}$ lies in $(-\varepsilon,\varepsilon)$, so that the inequality above implies $\sum_{i>N}f(x_{i})\leq\tfrac{1}{2}M\sum_{i>N}x_{i}^{2}\leq\tfrac{1}{2}M\lVert x\rVert_{\ell^{2}}^{2}$. For $\mathcal{C}_{1}$, $\mathcal{C}_{2}\in\mathcal{E}$ we have $R(\mathcal{C}_{2}\|\mathcal{C}_{1})\in L_{2}(\mathcal{H})$ and its eigenvalue sequence is square summable. Hence $\mathcal{L}_{f}<\infty$ by the definition of $\mathcal{L}_{f}$ in (16a). This proves item (i). For item (ii), we note $f(\eta(0))=f(0)=0$. Furthermore, we compute $\eta^{\prime}(x)=-(1+x)^{-2}$ and, by the fact that $f\in\mathscr{F}$, $x(f\circ\eta)^{\prime}(x)=\frac{1}{1+x}\frac{-x}{1+x}f^{\prime}(\frac{-x}{1+x})>0$ for $x\not=0$. By the assumption on $f$, $\lim_{x\rightarrow\infty}f(\eta(x))=\lim_{x\rightarrow-1}f(x)=\infty$. Finally, $\eta$ is smooth, so $\eta$ and $\eta^{\prime}$ are Lipschitz at 0. Therefore, $f^{\prime}\circ\eta$ is Lipschitz at 0 as the composition of Lipschitz functions at 0, and $(f\circ\eta)^{\prime}=(f^{\prime}\circ\eta)\eta^{\prime}$ is Lipschitz at 0 as the product of two Lipschitz functions at 0. ∎
@@
Hence $xf^{\prime}_{\textup{Ren},\rho}(x)>0$ for $x\not=0$. Furthermore, $f_{\textup{KL}}(0)=0=f_{\textup{Ren},\rho}(0)$ and $\lim_{x\rightarrow\infty}f_{\textup{KL}}(x)=\infty=\lim_{x\rightarrow\infty}f_{\textup{Ren},\rho}(x)$, so $f_{\textup{KL}},f_{\textup{Ren},\rho}\in\mathscr{F}$.
The first equations in items (i) and (ii) follow from Theorem 3.8. With $(\lambda_{i})_{i\in\mathbb{N}}$ the eigenvalues of $R(\mathcal{C}_{2}\|\mathcal{C}_{1})\in L_{2}(\mathcal{H})$, it holds by (12) that
@@
By the spectral mapping theorem—see e.g. [27, Theorem VII.1(e)] for a version that does not assume that $\mathcal{H}$ is defined over the complex field $\mathbb{C}$—the eigenvalues of $I+R(\mathcal{C}_{2}\|\mathcal{C}_{1})$ are $(1+\lambda_{i})_{i}$, and the eigenvalues of $A_{\rho}\coloneqq\bigl{(}I+R(\mathcal{C}_{2}\|\mathcal{C}_{1})\bigr{)}^{\rho-1}\bigl{(}\rho I+(1-\rho)(I+R(\mathcal{C}_{2}\|\mathcal{C}_{1}))\bigr{)}$ are $(\gamma_{i})_{i}$, with $\gamma_{i}\coloneqq(1+\lambda_{i})^{\rho-1}(\rho+(1-\rho)(1+\lambda_{i}))$, $i\in\mathbb{N}$. The eigenvalues of $A_{\rho}-I$ are then $(\gamma_{i}-1)_{i}$ and by (11),
$\det(A_{\rho})=\det(I+(A_{\rho}-I))=\prod_{i}(1+(\gamma_{i}-1))=\exp\left(\sum_{i}\log(\gamma_{i})\right)=\exp\left(2\rho(1-\rho)\sum_{i}f_{\textup{Ren},\rho}(\lambda_{i})\right).$
This shows item (ii) holds. Since $\lim_{x\rightarrow-1}f_{\textup{KL}}(x)=\infty=\lim_{x\rightarrow-1}f_{\textup{Ren},\rho}(x)$, item (iii) now follows directly from items (i) and (ii) and Lemma 4.1(ii). ∎